OpenStudy (anonymous):

How to simplify the expression (-3c^3 d^4 e^6) ^2

Nnesha (nnesha):

hint: \[\huge\rm (x^a y^b)^m = x^{a \times m}y^{b \times m}\]

OpenStudy (welshfella):

use the rule for exponents - what Nnesha said

OpenStudy (anonymous):

oh okay so it would be \[-3 c^{6} d ^{8} e ^{12} ?\]

OpenStudy (anonymous):

cause you would multiply the 2 with all the rest right?

OpenStudy (welshfella):

doont forget to include the coefficients in your calculations

Nnesha (nnesha):

-3 is raising to the 2nd power as wll

OpenStudy (welshfella):

(-3)^2 = 9

Nnesha (nnesha):

-3 is same -3^1 \[\rm so (-3^1)^2= (-3)^{1 \times 2}\]

OpenStudy (anonymous):

well i mean for the coefficients, wouldn't they be the letters?

OpenStudy (anonymous):

so the answer is 9^2 c^6 d^8 e^12?

OpenStudy (anonymous):

Cause there are no numbers

Nnesha (nnesha):

that was an example \[\huge\rm (x^a y^b)^m = x^{a \times m}y^{b \times m}\] everything which is inside the parentheses raising to the m power

OpenStudy (anonymous):

If yes, would the final answer be \[9c^{6} d ^{8} e ^{6}\]

Nnesha (nnesha):

\[\huge\rm (e^6)^2 = e^{??}\]

OpenStudy (anonymous):

oh i meant e^12

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

okay thank you

Nnesha (nnesha):

np